# OmniLens: Towards Universal Lens Aberration Correction via LensLib-to-Specific Domain Adaptation

### v2

@@ 4 Experiments and Results
AODLib
-LensNet [11]
FOV-KPN 4.01M 23.84 0.750 0.2466 41.09 63.54
@@
AODLib
-EAOD (Ours)
FOV-KPN 4.01M 25.62 0.785 0.1902 27.16 70.21
@@
Figure 5: Visual comparison with different LensLibs under 33 representative test lenses, which cover different surface types and levels of aberrations. The optical paths of the $3$ lenses are shown on the top of the figure. For each LensLib solution, we present its visual CAC results on an ISO12233 chart produced by FeMaSR which delivers the best perceptual performance.
4.1 Implementation Details
Evaluation Protocol. We manually design $6$ different specific lenses as the test samples, including $4$ spherical lenses (S-1P-I, S-1P-II, S-2P, and S-3P) and $2$ aspherical lenses (AS-1P and AS-3P) of different piece numbers. A synthetic benchmark of simulated images under each lens is then set up based on imaging simulation, where the referenced metrics, i.e., PSNR, SSIM [52], LPIPS [61], and FID [19] are employed for comprehensive evaluation. We also normalize and weight these metrics following [35], constructing an Overall Score (OS) to intuitively evaluate the performance of each solution (detailed in supplement). In addition, we fabricate two of the single spherical lenses to capture test images for constructing a real-world benchmark.
Data Preparation. The Flickr2K  [50] is chosen to generate $D_{S}$. In terms of synthetic evaluation, to avoid information leakage, target-specific training data $\mathcal{D}^{L_{j}}_{T}$ and test data are prepared with another dataset DIV2K to simulate the real-world captured images, where the ground-truth is unavailable during DA training. For real-world evaluation, we capture $50$ images as the test set, and another $25$ images as $\mathcal{D}^{L_{j}}_{T}$, with each aforementioned fabricated lens.
Model Training. We select FOV-KPN [4] (CNN and spatially-varying design), SwinIR [34] (Transformer), and FeMaSR [2] (CNN, Transformer, and generative model) as the representative networks in CAC to evaluate our OmniLens framework, all trained with default learning rates and loss functions for $200K$ iterations on $D_{S}$ to obtain the pre-training model. While for DA training, we use $L_{DA}$ with a window size of $15$ for DC image calculation to fine-tune the pre-trained model for $10K$ iterations with a learning rate of 1e61e{-}6, where $\lambda_{S}$ and $\lambda_{T}$ are set to $0.01$ and $1$ respectively, which is consistent across all $3$ networks.
More details can be found in the supplemental document.
4.2 Comparison with Different LensLibs
As shown in Tab. 1, we first verify the effectiveness of AODLib built by EAOD via comparison with the other two LensLib pipelines, i.e., the ZEBASELib and the ZernikeLib introduced in Sec. 2, whose construction process is consistent with [14] and [23] respectively. Additionally, the deep-learning-based AOD method LensNet [11] is also applied to construct a competing AODLib. The universal models with different architectures trained on our AODLib-EAOD outperform those under other LensLibs by a large margin, e.g., achieving the best OS, and bringing average gains of $1.78dB{\sim}1.93dB$ in PSNR and $0.084{\sim}{-}0.013$ in LPIPS. It can also be seen from the visual results in Fig. 5 that the CAC results of other pipelines still suffer from optical degradation under severe aberrations, where those of our method are sharp and clear. These evidences prove that the proposed EAOD method can build a more convincing LensLib for developing a robust universal CAC model generalizing well to most real-world lenses.
@@ 4.2 Comparison with Different LensLibs
Solution Network
Training Cost*
(Imgs/Iters)
Average Performance
@@
OmniLens
Zero-Shot
FOV-KPN 0/0K 25.62 0.785 0.1902 27.16 70.21
@@ 4.3 Evaluation of OmniLens
In lens-descriptions-unknown cases, we compare our pre-trained model and DA model in OmniLens with the state-of-the-art blind deconvolution method Eboli et al. [12]. For cases with known lens descriptions, lens-specific models under the selected $3$ networks are selected as baselines to evaluate OmniLens pre-training.
@@
Synthetic Benchmark. Tab. 2 shows the average performance of competing methods in the synthetic benchmark, where Fig. 6 also provides detailed overall results across each test lens. In the lens-descriptions-unknown situation, the proposed OmniLens significantly outperforms traditional blind deconvolution method. Moreover, DA effectively improves the pre-trained model by $0.22dB$ in PSNR and 0.0096-0.0096 in LPIPS. The improvements are observed across most test cases, which are more obvious under lenses with severe aberrations, e.g., S-1P-I, S-1P-II, and AS-1P. Then, with known lens descriptions, OmniLens pre-training brings improvements to the lens-specific models under both few-shot ($1.81dB$ in PSNR and 0.0527-0.0527 in LPIPS) and full training ($0.35dB$ in PSNR and 0.0088-0.0088 in LPIPS) settings. Notably, even the few-shot specific model with OmniLens outperforms the full-trained one, illustrating OmniLens’s potential to serve as a CAC pre-training foundation.
Figure 6: Evaluation of OmniLens across all test lenses. We show the OS results produced by FeMaSR with best performance.
@@
Real-world Benchmark. Fig. 7 presents some of the results (top $4$ solutions in Tab. 2) on real-world benchmark. Due to the synthetic-to-real domain gap, the results of lens-specific solution (③) reveal severe artifacts and unresolved optical degradation, while our OmniLens pre-training (① and ④) can effectively suppress the artifacts. Moreover, OmniLens-DA (②) shows great advantages in this case, effectively handling the optical degradation (image blur and purple fringes) without introducing visually unpleasant artifacts. Benefiting from the regularization of $\mathcal{L}_{CAC}$ on AODLib, the DA method can work in sky regions where DCP fails. The results of the user study are highly consistent with the visual results, where the results of OmniLens-based solutions are preferred by the majority of subjects.
@@ 4.4 Ablation Study
Ablations on EAOD. We explore the effectiveness of the components in EAOD as shown in Tab. 4. Fixing $S_{p}$, experiments 1-4 illustrate the significance of the evolution mechanism, physical constraints, and global optimization for guaranteeing AODLib’s wide and realistic coverage to develop a robust universal model. Then, in terms of the implementation of the hybrid optimization strategy, we suggest the application of SA for global optimization and ADAM for local optimization to achieve better results (experiments 3, 5, 6). Moreover, experiments 3 and 7-10 reveal that each specification contributes to building a superior AODLib, where the absence of any one will limit the diversity of lens samples, resulting in a decline in model performance.

### v3

@@ 4 Experiments and Results
AODLib -LensNet [cote2021deep] FOV-KPN 4.01M 23.84 0.750 0.2466 41.09 63.54
@@
AODLib -EAOD (Ours) FOV-KPN 4.01M 25.62 0.785 0.1902 27.16 70.21
@@
Figure 5: Visual comparison with different LensLibs under 33 representative test lenses, which cover different surface types and levels of aberrations. The optical paths of the $3$ lenses are shown at the top of the figure. For each LensLib solution, we present its visual CAC results on an ISO12233 chart produced by FeMaSR which delivers the best perceptual performance.
@@
Solution Network Training Cost* (Imgs/Iters) Average Performance
@@
OmniLens Zero-Shot FOV-KPN 0/0K 25.62 0.785 0.1902 27.16 70.21
@@
Figure 6: Evaluation of OmniLens across all test lenses. We show the OS results produced by FeMaSR with the best performance.
4.1 Implementation Details
Evaluation Protocol. We manually design $6$ different specific lenses as the test samples, including $4$ spherical lenses (S-1P-I, S-1P-II, S-2P, and S-3P) and $2$ aspherical lenses (AS-1P and AS-3P) of different piece numbers. A synthetic benchmark of simulated images under each lens is then set up based on imaging simulation, where the referenced metrics, i.e., PSNR, SSIM [wang2004image], LPIPS [zhang2018unreasonable], and FID [heusel2017gans] are employed for comprehensive evaluation. To intuitively evaluate the performance of each solution, we normalize and weight these metrics following [liang2024ntire], constructing an Overall Score (OS):
$\begin{split}&OS=20\times\frac{PSNR}{50}+15\times\frac{SSIM-0.5}{0.5}\\ &+20\times\frac{1-LPIPS}{0.4}+15\times\frac{100-FID}{100}.\end{split}$ (10)
In addition, we fabricate two of the single spherical lenses to capture test images for constructing a real-world benchmark. A commercial lens: $CAYE_{\circledR}50mm$, under different F-numbers, is also applied for further evaluation.
Data Preparation. The Flickr2K  [timofte2017ntire] is chosen to generate $D_{S}$. In terms of synthetic evaluation, to avoid information leakage, target-specific training data $\mathcal{D}^{L_{j}}_{T}$ and test data are prepared with another dataset DIV2K to simulate the real-world captured images, where the ground-truth is unavailable during DA training. For real-world evaluation, we capture $50$ images as the test set, and another $25$ images as $\mathcal{D}^{L_{j}}_{T}$, with each aforementioned fabricated lens.
Model Training. We select FOV-KPN [chen2021extreme_quality] (CNN and spatially-varying design), SwinIR [liang2021swinir] (Transformer), and FeMaSR [chen2022real] (CNN, Transformer, and generative model) as the representative networks in CAC to evaluate our OmniLens framework, all trained with default learning rates and loss functions for $200K$ iterations on $D_{S}$ to obtain the pre-training model. While for DA training, we use $L_{DA}$ with a window size of $15$ for DC image calculation to fine-tune the pre-trained model for $10K$ iterations with a learning rate of 1e61e{-}6, where $\lambda_{S}$ and $\lambda_{T}$ are set to $0.01$ and $1$ respectively, which is consistent across all $3$ networks.
More details can be found in the supplemental document.
4.2 Comparison with Different LensLibs
As shown in Tab. 1, we first verify the effectiveness of AODLib built by EAOD via comparison with the other two LensLib pipelines, i.e., the ZEBASELib and the ZernikeLib introduced in Sec. 2, whose construction process is consistent with [gong2024physics] and [jiang2024computational] respectively. Additionally, the deep-learning-based AOD method LensNet [cote2021deep] is also applied to construct a competing AODLib. The universal models with different architectures trained on our AODLib-EAOD outperform those under other LensLibs by a large margin, e.g., achieving the best OS, and bringing average gains of $1.78dB{\sim}1.93dB$ in PSNR and $0.084{\sim}{-}0.013$ in LPIPS. It can also be seen from the visual results in Fig. 5 that the CAC results of other pipelines still suffer from optical degradation under severe aberrations, where those of our method are sharp and clear. These evidences prove that the proposed EAOD method can build a more convincing LensLib for developing a robust universal CAC model generalizing well to most real-world lenses.
@@ 4.3 Evaluation of OmniLens
In lens-descriptions-unknown cases, we compare our pre-trained model and DA model in OmniLens with the state-of-the-art blind deconvolution method Eboli et al. [eboli2022fast]. For cases with known lens descriptions, lens-specific models under the selected $3$ networks are selected as baselines to evaluate OmniLens pre-training.
@@
Synthetic Benchmark. Tab. 2 shows the average performance of competing methods in the synthetic benchmark, where Fig. 6 also provides detailed overall results across each test lens. In the lens-descriptions-unknown situation, the proposed OmniLens significantly outperforms the traditional blind deconvolution method. Moreover, DA effectively improves the pre-trained model by $0.22dB$ in PSNR and 0.0096-0.0096 in LPIPS. The improvements are observed across most test cases, which are more obvious under lenses with severe aberrations, e.g., S-1P-I, S-1P-II, and AS-1P. Then, with known lens descriptions, OmniLens pre-training brings improvements to the lens-specific models under both few-shot ($1.81dB$ in PSNR and 0.0527-0.0527 in LPIPS) and full training ($0.35dB$ in PSNR and 0.0088-0.0088 in LPIPS) settings. Notably, even the few-shot specific model with OmniLens outperforms the full-trained one, illustrating OmniLens’s potential to serve as a CAC pre-training foundation.
Figure 8: Application case of OmniLens on a commercial lens with unknown lens descriptions. Results under different F-numbers with diverse aberration behaviors are provided for a comprehensive evaluation.
@@
Real-world Benchmark. Fig. 7 presents some of the results (top $4$ solutions in Tab. 2) on the real-world benchmark. Due to the synthetic-to-real domain gap, the results of the lens-specific solution (③) reveal severe artifacts and unresolved optical degradation, while our OmniLens pre-training (① and ④) can effectively suppress the artifacts. Moreover, OmniLens-DA (②) shows great advantages in this case, effectively handling the optical degradation (image blur and purple fringes) without introducing visually unpleasant artifacts. Benefiting from the regularization of $\mathcal{L}_{CAC}$ on AODLib, the DA method can work in sky regions where DCP fails. The results of the user study are highly consistent with the visual results, where the results of OmniLens-based solutions are preferred by the majority of subjects.
Figure 9: Detailed analysis of DCP in optical degradation. (a): The grayscale histograms for the DC images of the $6$ test lenses. We mark the exact number of zero pixels corresponding to each lens aberration. (b): For each lens’s DC images, we report the proportion of pixels with grayscale values less than “5” relative to the total number of pixels. We also provide the same statistics for the clear images as a reference. (c): Visual results of DC images under different severities of aberrations. (d): OS improvements brought to the pre-trained model by the DCP-based DA under different lenses.
Evaluation on the Commercial Lens. We further provide real-world results under a commercial lens: $CAYE_{\circledR}50mm$ in Fig. 8, where different apertures are applied for diverse F-number settings. Given that such a commercial lens lacks specific lens descriptions, we compare only the blind paradigm, namely the SOTA blind deconvolution method Eboli et al. [eboli2022fast] and our OmniLens-Zero-shot and OmniLens-DA. The results show the OmniLens pre-trained model effectively handles aberrations of such a commercial lens with an unknown lens description under large F-numbers. Meanwhile, the DA method further enhances CAC results, especially in severe aberrations with a small F-number. These cases with severe aberrations cannot be addressed by the traditional deconvolution method Eboli et al., indicating the favorable generalization of the OmniLens framework.
4.4 Evaluation of DCP in Optical Degradation
To analyze the DCP properties of optical degradation under different severities of aberrations, the $6$ test lenses are first divided into “weak”, “medium”, and “strong” aberration levels (detailed in the supplemental document). Using the same settings as the DCP analysis in Sec 3.5, we conduct a statistical analysis of the DC images for all $6$ test lenses, as shown in Fig 9. It can be observed that aberrations at all levels exhibit DCP, which is reflected in dark channel images with fewer zero pixels (Fig. 9 (a)), a smaller proportion of pixels with small gray values (Fig. 9 (b)), and “brighter” DC images (Fig. 9 (c)), illustrating the university of DCP in optical degradation. Meanwhile, the more severe aberrations reveal more obvious DCP properties. The experimental results in Fig. 9 (d) where more significant improvements in OS brought by DCP-based DA for “strong” aberrations further validate this.
@@ 4.5 Ablation Study
Ablations on EAOD. We explore the effectiveness of the components in EAOD as shown in Tab. 4. Fixing $S_{p}$, experiments 1-4 illustrate the significance of the evolution mechanism, physical constraints, and global optimization for guaranteeing AODLib’s wide and realistic coverage to develop a robust universal model. Then, in terms of the implementation of the hybrid optimization strategy, we suggest the application of SA for global optimization and ADAM for local optimization to achieve better results (experiments 3, 5, 6). Beyond performance, compared with the other LensLibs in Table 1, any combination of global and local optimization algorithms can produce a relatively high‑quality LensLib, indicating the advantage of the evolution mechanism in LensLib generation. Moreover, the local optimization algorithm, which determines the output multiple local optima, reveals a greater overall impact. In addition, experiments 3 and 7-10 reveal that each specification contributes to building a superior AODLib, where the absence of any one will limit the diversity of lens samples, resulting in a decline in model performance.
@@ 5.1 Conclusion
5.2 Discussion and Future Work
Enriching the LensLib. Considering the complexity of optimization, the EAOD currently only accommodates spherical lens structures, leaving room for improvements in the coverage of AODLib. Introducing additional degrees of freedom in optimization, such as aspheric and diffractive elements, could yield design outcomes with richer aberration behaviors, which can hardly be achieved by spherical lenses. Additionally, simulating manufacturing tolerances and assembly errors on optimized lens samples is highly valuable. These factors introduce stochastic, unknown aberration behaviors, making them a natural application case for blind aberration correction. Future work can center around incorporating more design specifications and perturbing LensLib samples within tolerance bands to diversify aberration distributions, thereby broadening dataset coverage and strengthening generalization for the pre-trained model.
Designs for Model Paradigm. The rich aberration behaviors of AODLib have not been fully leveraged during the training of our pre-trained model, as the model is merely fed with data from lenses with various aberrations in a random manner. We believe that appropriately classifying and annotating the aberration behaviors in AODLib, along with incorporating suitable prompts during model training, could enhance the generalization capability of the pre-trained models. In this context, we are particularly interested in utilizing prompt learning techniques in recent works of all-in-one image restoration.
Compensation for DCP Regularization. While DCP is promising as an unsupervised constraint for CAC, training instability and darker outputs indicate room for improvement. A practical path is to augment DCP with auxiliary low-level vision priors, as commonly done in image dehazing [chen2021psd, li2019semi]. For example, the bright channel prior can compensate for DCP failures in sky and specular regions, and the gradient sparsity prior, which is often expressed as Total Variation Loss (TV Loss), can further suppress noise and artifacts brought by the unsupervised DCP constraint.